in 10 years, william will be twice as old as he was 7 years ago. what is williams age now?
Answer by ankor@dixie-net.com(22740)

let a = his present age
:
in 10 years, william will be twice as old as he was 7 years ago.
a + 10 = 2(a-7)
a + 10 = 2a - 14
10 + 14 = 2a - a
a = 24 is williams age now?